OpenStudy (anonymous):

|4x-2|=26 there are two answers , please show work.

OpenStudy (saifoo.khan):

|4x-2|=26 4x - 2 = 26 AND 4x - 2 = -26

OpenStudy (saifoo.khan):

Solve both seperately.

OpenStudy (saifoo.khan):

separately